OpenStudy (anonymous):

the difference between the length of a hypotenuse and the length of the next longest side of a right triangle is 3 cm. the difference between the lengths of the two perpendicular sides is 3cm. find the three side lengths

OpenStudy (anonymous):

would someone please explain how i answer the question?

Directrix (directrix):

I think I know how to do this but I have to help someone finish a problem. Let h be the hypotenuse. Then, H = 3 + L-one where L-one is the longer leg. Let L-two = L-one minus 3 because L-two is the shorter of the two legs. The equation (Pythagorean Theorem) would be: h^2 =( Leg1)^2 + (Leg2)^2 (3 + L1)^2 = ( Leg1)^2 + (Leg2)^2 (3 + L1)^2 = ( Leg1)^2 + (3 + L1)^2 Now, solve for Leg One as denoted by L1. First, see if you agree with the equation.

Directrix (directrix):

Do you understand the equation?

OpenStudy (anonymous):

so shouldn't it be (3+L1)^2 = (leg 1)^2 + (L1 - 3)^2 ?? :S

OpenStudy (campbell_st):

 then its x^2= (x - 6)^2 + (x - 3)^2 x^2 = x^2 - 12x + 36 + x^2 - 6x + 9 collect like terms and simplify x^2 - 18x + 45 = 0 (x - 15)(x - 3) = 0 x = 15, 3 the solution of x = 3 can't be used as the side lengths would be x=3, 0 and -3 use x = 15 Hypotenuse = 15, shorter side = 12 shortest side = 9 ( 3,4,5 triad)